OpenStudy (anonymous):

Convert 5i-3j to polar coordinates. (write 4 forms for this point.)

OpenStudy (anonymous):

Hi! Have you tried drawing a diagram?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Polar_coordinate_system r=(25+9)^0.5=34^0.5 \[\psi=\arctan(-3/5)\], \[(\psi,r)\] is the answer